OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 3x + 5y = 3 x + 2y = 0 I keep getting the wrong answer. Can anybody show me how to solve this?

jhonyy9 (jhonyy9):

x=-2y -6y+5y=3 -y=3 y=-3 x=6

OpenStudy (anonymous):

Thanks!